OpenStudy (anonymous):

How to calculate or simplify the the integral of ((18(ln(x))^5)/x)dx ?

OpenStudy (adamaero):

Then, the property of ln() is that if it's raised to anything, you can take that out in front: ln(x)^2 = 2ln(x) Take it from there...

OpenStudy (amistre64):

by parts maybe? let u = ln(x) and dv = 1

OpenStudy (anonymous):

\[\int\limits \frac{ 18\left( \ln \left( x \right) \right)^5 }{ x }dx\]

OpenStudy (anonymous):

like this?

OpenStudy (amistre64):

theres an x under there eh .... then just a u sub seems sufficient

OpenStudy (anonymous):

if so, let u = ln(x), du = (1/x) dx

OpenStudy (anonymous):

Yes pgpilot326 ! thank you

OpenStudy (anonymous):

And how do you integrate an ln function? I can't find a clear definition of it in my book.

OpenStudy (amistre64):

\[ln^5(x)\ne ln(x^5)\] so that exponent property doesnt fit

OpenStudy (amistre64):

the key to this is in knowing how to take the derivative of the ln function

OpenStudy (anonymous):

in this one you don't have to integrate ln(x)

OpenStudy (amistre64):

when that becomes clear, it can be seen that the chainrule was applied and it cleans up with the u subsititution

OpenStudy (anonymous):

I think I got it. Does it become 18(u)^5 du?

OpenStudy (amistre64):

it does :)

OpenStudy (anonymous):

Ok thank you all very much!

OpenStudy (anonymous):

\[\int\limits \frac{ 18\left( \ln \left( x \right) \right)^5 }{ x }dx=18\int\limits u^5 du=\frac{ 18u^6 }{ 6 }=3u^6 = 3\left( \ln \left( x \right) \right)^6\]

OpenStudy (amistre64):

waiting for someone to shime in with the +C

OpenStudy (anonymous):

i'm singing it... 1 octave above middle C